### v1

# LongFaith: Enhancing Long-Context Reasoning in LLMs with Faithful Synthetic Data

Abstract
Despite the growing development of long-context large language models (LLMs), data-centric approaches relying on synthetic data have been hindered by issues related to faithfulness, which limit their effectiveness in enhancing model performance on tasks such as long-context reasoning and question answering (QA). These challenges are often exacerbated by misinformation caused by lack of verification, reasoning without attribution, and potential knowledge conflicts. We propose LongFaith, a novel pipeline for synthesizing faithful long-context reasoning instruction datasets. By integrating ground truth and citation-based reasoning prompts, we eliminate distractions and improve the accuracy of reasoning chains, thus mitigating the need for costly verification processes. We open-source two synthesized datasets—LongFaith-SFT and LongFaith-PO—which systematically address multiple dimensions of faithfulness, including verified reasoning, attribution, and contextual grounding. Extensive experiments on multi-hop reasoning datasets and LongBench demonstrate that models fine-tuned on these datasets significantly improve performance. Our ablation studies highlight the scalability and adaptability of the LongFaith pipeline, showcasing its broad applicability in developing long-context LLMs.
LongFaith: Enhancing Long-Context Reasoning in LLMs with Faithful Synthetic Data
Cehao Yang1,2111Equal contribution., Xueyuan Lin1,2111Equal contribution., Chengjin Xu1111Equal contribution., Xuhui Jiang1, Shengjie Ma1, Aofan Liu1, Hui Xiong2222Corresponding authors., Jian Guo1222Corresponding authors., 1IDEA Research, International Digital Economy Academy 2Artificial Intelligence Thrust, Hong Kong University of Science and Technology (Guangzhou) {yangcehao,linxueyuan,xuchengjin,jiangxuhui,mashengjie,liuaofan,guojian}@idea.edu.cn {cyang289,xlin058}@connect.hkust-gz.edu.cn, xionghui@ust.hk
33footnotetext: Our code implementation and datasets can be accessed at https://github.com/IDEA-FinAI/LongFaith.
1 Introduction
Long-context processing ability has emerged as a significant challenge for large language models (LLMs) Shi et al. (2023); Liu et al. (2024); Wu et al. (2024); Levy et al. (2024), especially arises when models process extensive textual information, making it hard to recognize relevant evidence and address downstream tasks such as question answering (QA), summarization, and complex reasoning Bai et al. (2023, 2024b); Zhang et al. (2024d); Hsieh et al. (2024); Yen et al. (2024). A variety of model-centric methods have been proposed to extend the length of context windows in LLMs Chen et al. (2023a, b); Peng et al. (2023); Han et al. (2024); Ding et al. (2024). Additionally, many data-centric methods, such as synthesizing long-context understanding instructions for fine-tuning, have gained attention for enhancing models’ ability to handle and utilize extended contexts Xiong et al. (2023); An et al. (2024); Fu et al. (2024); Bai et al. (2024a); Li et al. (2024b); Gao et al. (2024); Chen et al. (2024); Zhang et al. (2024c); Li et al. (2024a).
Figure 1: A brief introduction of LongFaith. Synthesized long-context reasoning instruction sets and preference datasets are fed into the post-training stage of downstream LLMs.
Figure 2: Overview of LongFaith pipeline for synthesizing faithful long-context reasoning instruction and preference datasets. Comparing generated reasoning chains with misinformation, lack of attribution, and knowledge conflicts, LongFaith generates ground truth guidance prompting by chain-of-citation to build LongFaith-SFT. Fine-grained faithfulness is modeled by optimization on our preference datasets LongFaith-PO.
Despite the improvements in downstream QA performance enabled by synthetic long-context reasoning instructions, concerns remain regarding the faithfulness of such generated data. Specifically: (1) Misinformation due to lack of verification: existing methods often generate QA pairs without rigorous rule-based verification. For instance, Chen et al. (2024); Zhang et al. (2024c); Li et al. (2024a) directly synthesize QA pairs using LLMs while bypassing verification, whereas Zhang et al. (2024c) relies on AI-generated feedback in soft dimensions rather than human annotation. (2) Reasoning without attribution: prompting LLMs to generate responses with citation, such as using Chain-of-Citation (CoC) prompting Li et al. (2023, 2024b); Fierro et al. (2024); Huang et al. (2024); Berchansky et al. (2024); Gao et al. (2023) can enhance the credibility and interpretability of model outputs under long-context QA tasks Gao et al. (2023); Zhang et al. (2024b), yet most prior works ignore to incorporate this technique during their synthesis of training instruction pairs. (3) Potential knowledge conflicts: some approaches Bai et al. (2024a); Zhang et al. (2024c); Chen et al. (2024) over-rely on the Self-Instruct technique Wang et al. (2022) to generate QA pairs, encouraging models to rely on parametric knowledge rather than grounding reasoning in explicit contextual evidence Xu et al. (2024). Additionally,  Zhang et al. (2024c) feeds the query and response to a short-context reward model ignoring the long context to score, purely relying on the parametric knowledge inside LLM. These limitations underscore the necessity for a more robust pipeline that ensures the faithfulness of long-context instructions synthesis.
We propose LongFaith, a novel pipeline for synthesizing faithful long-context reasoning instruction datasets. We incorporate ground truth directly into the prompt for synthesizing long-context reasoning chains, which comprise supporting facts and the correct answer, and prompt LLMs to reason with attributions. This method ensures the faithfulness of synthesized reasoning chains without requiring costly verification by a curated rule-based evaluator, LLM-as-a-judge Gu et al. (2024) or human annotator. We open-source LongFaith-SFT, synthesized under the guidance of ground truth and CoC prompting. We leverage the faithful long-context reasoning chains with attributions for training, leading to performance improvements after fine-tuning Llama-3.1-8B-Instruct. Additionally, we synthesize preference datasets by sampling preference pairs around fine-grained faithfulness: (1) encouraging model to reason with attributions; (2) encouraging model to learn on verified reasoning chains; and (3) encouraging model to reason with contextual documents grounded. We open-source LongFaith-PO, synthesized by various LLMs in different sizes, which integrates all three faithfulness dimensions for preference optimization. We leverage these faithful preferred instruction pairs for training Llama-3.1-8B-Instruct, achieving performance improvements on the multi-hop reasoning dataset and LongBench Bai et al. (2023).
Our main contributions are as follows: (1) We introduce LongFaith, a novel pipeline for synthesizing faithful long-context reasoning instruction data. (2) We open-source LongFaith-SFT and LongFaith-PO, two synthesized datasets that are systematically designed considering multiple dimensions of faithfulness. (3) We conduct extensive experiments on two types of datasets (comprising eight sub-tasks) to show that models trained on LongFaith datasets can improve in long-context reasoning and QA tasks. (4) Our ablation studies illustrate the scaling potential and adaptability of LongFaith pipeline, underscoring its broad applicability in the development of long-context LLMs.
2 Related Work
Long-Context Utilization.
Amounts of studies focus on enhancing LLMs to better utilize long-context information. Model-centric approaches, for instance, optimizations on attention mechanism aim to capture specific sequential features Beltagy et al. (2020); Ding et al. (2023); Chen et al. (2023b); Han et al. (2024), while positional interpolation techniques are utilized to scale positional encoding while ensuring valid index ranges Zhu et al. (2023); Chen et al. (2023a); Ding et al. (2024); Peng et al. (2023); Beltagy et al. (2020). In addition, data-driven approaches also gain popularity, emphasizing high-quality data synthesis for fine-tuning to improve LLMs’ long-context processing capabilities. For example, Xiong et al. (2023); Gao et al. (2024) employ long-sequence continuous pre-training on foundation models, while Fu et al. (2024) explores the impact of pre-training data composition and balance. Additionally, works on SFT with synthetic instructions An et al. (2024); Bai et al. (2024a); Li et al. (2024b); Chen et al. (2024) not only consider long-context understanding but also strengthen multi-hop reasoning capabilities. Lastly, preference optimization approaches Zhang et al. (2024c); Li et al. (2024a) generate fine-grained pairwise preference instruction sets and incorporate training techniques Rafailov et al. (2024); Hong et al. (2024). From the perspective of improving the faithfulness of synthetic data, our work effectively addresses the shortcomings of prior studies in this area.
Faithful Reasoning.
Hallucination in LLMs presents a major challenge in knowledge-intensive tasks Zhang et al. (2023); Huang et al. (2023). Recent work has focused on enhancing faithful reasoning, where the goal is to trace the LLM’s generated content back to reliable sources and ensure its factual grounding.  Berchansky et al. (2024); Li et al. (2023, 2024b) aim to improve the identification and verification of attributions by focusing on generating reasoning outputs that link claims to specific sources. Benchmarks such as  Gao et al. (2023); Yue et al. (2023) evaluate the quality of citations and highlight the limitations of current systems in providing citation support to ensure more reliable output. Additionally, integrating external knowledge sources has gained attention, which use retrieval-augmented generation (RAG) methods to facilitate deep and faithful reasoning Sun et al. (2023); Ma et al. (2024). Our LongFaith is motivated by previous work, towards faithful reasoning in long-context reasoning tasks.
3 LongFaith
In this section, we present an exposition of LongFaith pipeline. Specifically, we explain how it synthesizes LongFaith-SFT for supervised fine-tuning and LongFaith-PO for preference optimization from the perspective of faithfulness.
Synthesize Reasoning Chains with High Faithfulness.
Previous studies Bai et al. (2024a); Chen et al. (2023b, 2024); Zhang et al. (2024c); Li et al. (2024a) tend to directly distill synthesized long-context QA and reasoning instructions for training without filtering out incorrect information. These low-faithfulness synthesized data limit the performance improvements of the trained models. In response to this challenge, LongFaith integrates ground truth into the synthesized reasoning chains. For a sample $S$ from the training set $S=(Q,D,A,F)$, where $Q$ is the reasoning question, $D$ is the full document used for querying, $A$ is the correct answer, and $F$ represents the supporting facts where $F\in D$. We use the LLM $M_{\text{syn}}$ to synthesize the reasoning chain as follows:
$O_{c}=M_{\text{syn}}(P_{\text{coc}},Q,F,A)$ (1)
Here, $O_{c}$ represents the output of $M_{\text{syn}}$, which is a step-by-step reasoning chain. The prompt $P_{\text{coc}}$ utilizes a chain-of-citation Li et al. (2024b) prompting approach, requiring the model to reason with attribution (e.g., "Let’s reason step by step while citing the document using [1], [2], etc."). The prompt template is shown in Figure 7.
LongFaith-SFT Dataset.
Towards training a downstream LLM to reason with high faithfulness for a long-context QA task, we construct the dataset for supervised fine-tuning, where each instruction pair is built as follow:
$I_{\text{sft}}=\{\text{input}=(P_{\text{coc}},Q,D),\text{output}=O_{c}\}$ (2)
Synthesize Reasoning Chain with Questionable Faithfulness.
To model fine-grained preferences, we address three challenges that affect the faithfulness of synthesized instructions: (1) misinformation due to lack of verification, (2) reasoning without attribution, and (3) potential knowledge conflicts. We synthesize reasoning chains with questionable faithfulness, including reasoning chains with misinformation as follows:
$O_{m}=M_{\text{syn}}(P_{\text{coc}},Q,D)$ (3)
Since there is no ground truth to guide the synthesis, the output $O_{m}$ may contain errors in reasoning, as illustrated in Figure 8, where the model generates an incorrect answer of "1903" instead of the correct answer "1698". This hallucination is common in synthesized data from previous works unless rules or human experts are involved in filtering Li et al. (2024b). Next, we synthesize reasoning chains without attribution:
$O_{\text{cot}}=M_{\text{syn}}(P_{\text{cot}},Q,F,A)$ (4)
Here, the CoT Wei et al. (2022) prompting only requires the model to provide step-level reasoning, but as shown in Figure 9, reasoning without attribution not only loses interpretability and credibility Gao et al. (2023); Li et al. (2023), but our results in Tab. 4 (Sec. 4) also demonstrate that CoT prompting performs worse than CoC. Finally, we synthesize reasoning chains with potential knowledge conflicts:
$O_{\text{kc}}=M_{\text{syn}}(P_{\text{cot}},Q,A)$ (5)
Since no context is provided, the model relies solely on its parametric knowledge for reasoning, as shown in Figure 10, where the model states, "Panama was not colonized by the United Kingdom; Panama was colonized by Spain," based on internal parametric knowledge rather than the contextual documents. Previous studies Zhang et al. (2024c) using short-context reward models observes performance degradation by ignoring long-context materials, highlighting the limitation of knowledge conflicts in affecting LLM’s performance in long-context QA and reasoning tasks.
Synthesis of Reasoning Chains
Models Prompt w/ GT w/ Doc Output Size
Q-7B CoC 1 1-8K
Q-7B CoT 2 1-8K
Q-7B CoC 3 1-8K
Q-7B CoT 4 1-8K
L8,L70,G CoC 5 2K
L8,L70,G CoT 6 2K
L8,L70,G CoC 7 2K
L8,L70,G CoT 8 2K
Datasets for Supervised Fine-tuning
Name Models Instruction Output Size
LF-SFT Q-7B CoC 1 1-8K
w/o CoC Q-7B CoT 2 1-8K
w/o GT Q-7B CoC 3 1-8K
w/o Doc Q-7B CoC 4 1-8K
LF-SFT L8,L70,G CoC 5 2K
w/o CoC L8,L70,G CoT 6 2K
w/o GT L8,L70,G CoC 7 2K
w/o Doc L8,L70,G CoC 8 2K
Datasets for Preference Optimization
Name Models Instruction Chosen Rejected Size
w/ CoC Q-7B CoC 1 2 1-8K
w/ GT Q-7B CoC 1 3 1-8K
w/ Doc Q-7B CoC 1 4 1-8K
LF-PO Q-7B CoC 1 2,3,4 1-8K
w/ CoC L8,L70,G CoC 5 6 2K
w/ GT L8,L70,G CoC 5 7 2K
w/ Doc L8,L70,G CoC 5 8 2K
LF-PO L8,L70,G CoC 5 6,7,8 2K
Table 1: Statistics of synthesized datasets for SFT and PO. We first synthesize large-scale reasoning chains and then refactor them to datasets, where the second stage does not require llm inference. Q-7B means Qwen2.5-7B-Instruct, L8 means Llama-3.1-8B-Instruct, L70 means Llama-3.1-70B-Instruct and G means GPT-4o mini. GT means Ground Truth, CoC means chain-of-citation, Doc means contextual documents, and LF means LongFaith. 1-8K includes {1K, 2K, 4K, 8K}.
Llama-3.1-8B-Instruct MuSiQue 2Wiki HotpotQA Qasper(S) MFQA-En(S) MuSiQue(M) 2Wiki(M) HotpotQA(M)
F1 SubEM F1 SubEM F1 SubEM F1 SubEM F1 SubEM F1 SubEM F1 SubEM F1 SubEM
Zero-Shot Prompting
+ CoT 15.9 56.8 34.0 83.8 20.8 78.6 3.2 22.0 5.7 29.3 14.1 43.5 30.1 77.0 13.4 60.5
+ CoC 25.8 64.2 43.6 86.2 32.7 76.6 4.6 26.0 7.0 32.7 11.8 41.0 28.1 79.5 19.9 58.0
Superivised Fine-tuning
+ LongAlpaca 21.6 50.2 47.8 80.4 32.7 76.6 5.7 25.0 5.8 30.7 8.5 48.5 25.4 77.0 12.5 61.0
+ LongAlign 24.8 48.4 55.6 84.2 51.0 79.2 6.5 24.0 10.7 38.7 15.0 40.0 33.4 76.5 35.8 61.0
+ MuSiQue-Attribute 13.9 19.2 23.9 49.6 20.2 37.2 10.0 11.5 8.3 12.0 15.2 26.5 21.2 55.0 25.6 41.0
+ LongMIT 4.9 33.0 3.3 58.0 10.1 63.6 9.5 18.5 5.6 30.0 7.5 29.0 3.6 55.5 23.7 50.0
+ LongReward-SFT 6.2 48.4 23.3 80.0 15.6 74.2 2.6 22.5 0.5 34.0 1.1 43.0 6.6 71.5 8.9 54.0
+ SeaLong-SFT 31.3 64.6 55.8 89.2 59.4 83.0 14.5 26.0 18.6 31.3 24.1 59.5 34.1 84.5 37.3 69.0
+ LongFaith-SFT 56.8 62.8 73.8 85.6 70.5 80.8 36.9 29.5 47.0 32.0 50.1 56.5 63.9 82.0 53.1 68.0
Preference Optimization
+ LongReward-PO 3.3 46.0 14.3 76.6 8.9 71.2 1.6 21.0 0.1 32.7 0.0 37.5 4.4 67.0 3.3 53.0
+ SeaLong-PO 30.2 60.4 50.1 89.4 58.3 83.4 17.1 28.0 20.1 32.0 18.1 53.3 34.0 86.0 40.2 69.5
+ LongFaith-PO 60.5 66.4 68.0 85.0 65.4 81.2 38.1 30.5 46.7 32.0 50.2 52.0 73.7 83.5 55.6 67.5
Table 2: Main experiment results on three multi-hop reasoning test sets and five long-context QA test sets from LongBench. The best results are in bold and second-best are underlined. (S) means single-doc QA task and (M) means multiple-docs QA task in LongBench. LongFaith-SFT and LongFaith-PO are synthesized by GPT-4o mini both in 2K size. To ensure fairness, we sample first 2K examples from baseline datasets.
LongFaith-PO Dataset.
Towards training a downstream LLM to address three challenges above in long-context reasoning, we force the LLM to learn reasoning with high faithfulness while rejecting outputs of questionable faithfulness:
$\begin{split}I_{\text{po}}=\{\text{input}=(P_{\text{coc}},Q,D),\\ \text{chosen}=O_{c},\text{rejected}=O_{r}\}\end{split}$ (6)
where $O_{r}$ is a combination of ($O_{m}$, $O_{\text{cot}}$, $O_{\text{kc}}$), or a subset of them.
4 Experiments
Llama-3.1-8B-Instruct MuSiQue 2WikiMultiHopQA HotpotQA
Overall 2-Hop 3-Hop 4-Hop Overall 2-Hop 4-Hop Overall Bridge Comparison
Zero-Shot Prompting
+ CoT 11.0 7.5 16.2 12.0 29.0 22.3 54.3 17.4 17.5 17.0
+ CoC 19.0 16.1 22.7 20.7 39.2 31.4 68.6 30.4 28.6 38.6
Superivised Fine-tuning
+ LongFaith-SFT 40.6 44.1 37.7 35.9 55.4 51.1 71.4 53.6 57.0 37.5
    w/o CoC 40.2 41.7 39.6 37.0 51.8 48.9 62.9 52.0 56.6 30.7
    w/o GT 30.4 31.9 28.6 29.3 55.8 49.6 79.0 56.6 54.9 64.8
    w/o Doc 20.0 23.6 18.2 13.0 55.8 47.1 88.6 47.4 45.9 54.5
Preference Optimization
    w/ GT-PO 44.0 45.7 42.9 41.3 56.0 50.4 77.1 54.4 58.3 36.4
    w/ CoC-PO 43.6 44.5 44.8 39.1 53.2 48.6 70.5 56.2 59.2 42.0
    w/ Doc-PO 41.4 42.5 40.3 40.2 56.0 52.7 68.6 56.4 59.5 42.0
+ LongFaith-PO 46.6 47.2 48.1 42.4 59.0 55.9 70.5 58.6 59.7 53.4
Table 3: Main experiment results on three long-context multi-hop reasoning datasets using the Exact-Match(EM) metric. The best results are in bold. The training set has 2K samples for both SFT and PO, synthesized by Qwen2.5-7B-Instruct. Detail statistics of synthetics datasets are presented in Tab. 1.
4.1 Implementation Details
Following previous studies, we leverage the training set of MuSiQue Trivedi et al. (2022b), which is build on Wikipedia documents with supporting documents and correct answers. The officially retrieved 20 documents are provided and read only once in the input context in distractor setting. The statistics of training set is given in Tab. 5, covering 1K, 2K, 4K and 8K, where the balance of questions with different hops are considered. Following the pipeline we describe in Sec. 3, reasoning chains are samples to build LongFaith-SFT and LongFaith-PO. The statistics are presented in Tab. 1.
We conduct our experiments on a Linux server equipped with 4 A100-SXM4-40GB GPUs. For data synthesis of long-context reasoning instructions, we take Llama-3.1-8B-Instruct Dubey et al. (2024), Qwen2.5-7B-Instruct Yang et al. (2024), Llama-3.1-70B-Instruct and GPT-4o mini Hurst et al. (2024) as generators and prompt LLMs to synthesize reasoning chains with vLLM Kwon et al. (2023). We adopt the LoRA technique Hu et al. (2021) for fine-tuning and ORPO technique Hong et al. (2024) for preference optimization using the LLaMA-Factory framework Zheng et al. (2024) to train Llama-3.1-8B-Instruct. Hyperparameters of post-training are given in App. F.
Figure 3: Performance of Llama-3.1-8B-Instruct trained on different size of instructions synthesized by Qwen2.5-7B-Instruct from 1K to 8K, evaluated by EM and F1 metrics on multi-hop reasoning sets and LongBench.
4.2 Evaluation Setup
Following prior work Li et al. (2024b), we utilize three multi-hop reasoning datasets, including MuSiQue Trivedi et al. (2022b), 2WikiMultiHopQA Ho et al. (2020), and HotpotQA Yang et al. (2018), evaluating in distractor-setting, where the officially retrieved 10 or 20 documents are provided and read only once in the input context. We adopt the test sets sampled by Trivedi et al. (2022a), with 500 examples in each set. Furthermore, in line with previous studies Chen et al. (2024); Zhang et al. (2024c); Li et al. (2024a), we assess the performance on LongBench Bai et al. (2023), which includes two test sets for single-doc QA including Qasper Dasigi et al. (2021) and MultiFieldQA-EN Bai et al. (2023), as well as three test sets for multi-docs QA tasks including HotpotQA, 2WikiMultiHopQA, and MuSiQue. Notably, although there is an overlap in multi-hop reasoning tasks, the LongBench version further extends the lengths of document text. To apply CoC prompting, single document is split into 20 even paragraphs with order. The statistics of datasets are listed in Tab. 6.
To ensure fairness, we use Substring Exact-Match (SubEMYen et al. (2024); Li et al. (2024a) metric in main experiments, in case that models trained on baseline datasets are not good at instructions following to summarize the answer with "The answer is", and SubEM goes through the whole response to check whether the answer is in. Furthermore, following previous work Choi et al. (2018); Zhang et al. (2024a); Li et al. (2024b), we use EM metric and F1 scores for the trimmed part after "The answer is" for evaluation in main experiments and ablation studies. Comparing with LLM-as-a-Judge Bai et al. (2024a); Chen et al. (2024); Zhang et al. (2024c) using strong API models like GPT-4o, the rule-based metrics cost much lower.
4.3 Baselines
We compete LongFaith-SFT and LongFaith-PO with datasets proposed in previous studies, including LongAlpaca Chen et al. (2023b), LongAlign Bai et al. (2024a), MuSiQue-Attribute Li et al. (2024b), LongMIT Chen et al. (2024), LongReward-SFT Zhang et al. (2024c), SeaLong-SFT Li et al. (2024a) for supervised fine-tuning, and LongReward-PO and SeaLong-PO for preference optimization. All of them aim at enhancing the performance of LLMs on long-context understanding, reasoning, and QA tasks. To ensure fairness, we keep the training setting consistent with App. F and truncate the size of training samples to a maximum of 2K.
Figure 4: Scatter plot with a linear regression line fitting the relationship between QA - EM and Attribution - F1 metrics on three long-context multi-hop reasoning test sets. A point refers to the performance of a model trained with a specific size between 1K to 8K by SFT or PO.
4.4 Main Results
LongFaith Outperforms Previous Datasets.
Following previous work and to ensure a fair comparison, we evaluate the performance of LongFaith on multi-hop reasoning test sets Trivedi et al. (2022b); Ho et al. (2020); Yang et al. (2018) and LongBench Bai et al. (2023), comparing it against baseline methods, including zero-shot prompting with Llama-3.1-8B-Instructand models trained on synthetic datasets from previous works. As shown in Tab. 2, LongFaith outperforms baseline datasets on most test sets. The performance of the model trained on LongFaith-PO surpasses that trained on LongFaith-SFT. This aligns with our expectations: compared to directly using positive samples for supervised fine-tuning, incorporating rejected samples to provide more fine-grained faithfulness preferences for optimization leads to better improvements in long-context reasoning and QA capabilities. We observe that some synthetic instruction sets degrade performance compared to native Llama-3.1-8B-Instruct. This proves that datasets with questionable faithfulness are even harmful to long-context reasoning ability of LLMs.
LongFaith Arrives at the Correct Answer without Redundant Exploration.
We find that on 2WikiMultiHopQA, HotpotQA, and part of tasks in LongBench, SeaLong achieves a slight advantage in the SubEM metric against LongFaith, but fails in F1 scores. We investigate the length of response and present in Tab. 8. It turns out that the LLM trained on SeaLong conducts redundant exploration in response, producing more noisy content related to the answer, but actually arrives at a wrong answer, which means SubEM metric is easily to be "hacked". In contradiction, F1 scores requires to truncate the part after "The answer is", which demonstrates that a model trained on LongFaith datasets can arrive at the correct answer without redundant exploration and achieve a high score in a more strict metric. A case study is shown in Fig. 11 in Appendix.
Generalization.
Based on statistics from Tab. 7, the main experiment demonstrates that LongFaith uses instructions with shorter context as input compared to baseline methods, reducing training costs while generalizing to LongBench tasks that require processing an average of 24K-70K tokens as input. This further highlights the generalization ability of our pipeline.
Llama-3.1-8B-Instruct MuSiQue 2WikiMHQA HotpotQA LongBench(S) LongBench(M)
    + LongFaith EM F1 EM F1 EM F1 EM F1 EM F1
Superivised Fine-tuning
    w/ Llama-3.1-8B-Instruct 35.4 48.4 59.4 69.5 54.6 67.7 9.9 29.0 42.2 48.4
    w/ Qwen2.5-7B-Instruct 40.6 54.0 55.4 65.0 53.6 69.5 14.1 38.4 43.0 53.8
    w/ Llama-3.1-70B-Instruct 44.8 58.1 54.0 64.4 54.4 69.5 16.0 41.0 44.8 56.7
    w/ GPT-4o mini 41.6 56.8 64.6 73.8 55.4 70.5 16.9 42.0 47.2 55.7
Preference Optimization
    w/ Llama-3.1-8B-Instruct 41.2 53.2 57.4 67.0 55.6 68.5 14.7 36.9 44.0 55.3
    w/ Qwen2.5-7B-Instruct 46.6 59.2 59.0 67.4 58.6 72.8 15.4 35.3 44.8 55.6
    w/ Llama-3.1-70B-Instruct 50.4 63.2 52.8 62.7 57.2 71.0 16.4 40.0 48.3 59.4
    w/ GPT-4o mini 48.4 60.5 59.8 68.0 49.8 65.4 15.9 42.4 45.0 59.8
Table 4: Ablation study on various LLMs for synthesizing LongFaith-SFT and LongFaith-PO in the size of 2K. The base model for training and testing is Llama-3.1-8B-Instruct. (S) and (M) refer to Single-doc QA and Multi-docs QA in LongBench.
Figure 5: Visualization of F1 scores in Tab. 2.
4.5 Analysis
Exploration on Different Perspective of Faithfulness.
To validate the specific impact of different dimensions of faithfulness, we fine-tune models using negative samples as output and optimize using preference datasets that reject only a subset of negative samples. The statistics of the constructed datasets are shown in Tab. 1. Since each task in LongBench contains no more than 200 questions, performance evaluations can be prone to errors, so we chose to test on multi-hop reasoning datasets. Experimental results are shown in 3. The models trained with LongFaith-SFT and LongFaith-PO achieved high performance respectively in SFT and PO especially in F1 scores, as expected.
However, we note that in 4-Hop part of 2Wiki and comparison part of HotpotQA, LongFaith-SFT w/o CoC and w/o GT demonstrated better performance. Analysis reveals that for the question "Do both films, Cuban Colony and Prathyartha, have directors from the same country?", as the training set MuSiQue used specific entities as answers, the model responds "Both directors are from the same country, which is India. The answer is India.". Actually, the correct answer is "yes." Model trained on LongFaith w/o GT and LongFaith w/o Doc performed better with more exploration, but also lost overall performance due to hallucinations. Models trained on all PO datasets outperformed those trained using only positive samples for SFT, demonstrating the performance improvement brought by each fine-grained, credible preference. Finally, models trained on LongFaith-PO, which integrates three dimensions of faithfulness, achieved the best overall performance.
Scalability and Performance Gains.
We explore the scaling-up potential of LongFaith on multi-hop reasoning test sets and LongBench. As presented in Tab. 6, we train Llama-3.1-8B-Instruct using LongFaith-SFT and LongFaith-PO synthesized by Qwen2.5-7B-Instruct across four dataset sizes, ranging from 1K to 8K. According to the performance trend in Fig. 3, LongFaith generally exhibits scaling-up potential, indicating that expanding the training dataset can further enhance performance. Moreover, LongFaith-PO, which incorporates fine-grained preference optimization, demonstrates a more stable upward trend compared to LongFaith-SFT, particularly in LongBench tasks. This result validates the robustness of the LongFaith pipeline.
Attribution-Based Reasoning Leads to Higher Performance.
Utilizing CoC prompting for reasoning with attributions not only outperforms CoT in performance, as it presents in Tab. 3, but also provides greater interpretability and faithfulness as shown in Fig. 2. We use Attribution F1 as a metric to quantify the model’s attribution capability using annotated supporting facts. Under CoC prompting, we analyze the references within reasoning chains, matching them against supporting facts like [1], [2], etc., and compute F1 scores based on recall and precision. We evaluate the attribution capability and overall performance of Llama-3.1-8B-Instructtrained on LongFaith-SFT and LongFaith-PO across four sizes and visualize the results in a scatter plot. The findings in 4 demonstrate a strong positive correlation between attribution capability and model performance, further validating the effectiveness of the LongFaith pipeline.
Impact of LLM Selection for Synthesis.
We experimented with different LLMs for synthesis, including smaller open-source LLMs such as Llama-3.1-8B-Instruct, Qwen2.5-7B-Instruct, and larger open-source models like Llama-3.1-70B-Instruct, as well as a closed-source API model, GPT-4o mini, to synthesize LongFaith-SFT and LongFaith-PO for training Llama-3.1-8B-Instruct. The performance test results are presented in Tab.4. Using stronger closed-source API models to synthesize LongFaith-SFT led to a stronger performance boost, which aligns with intuition and previous work Chen et al. (2024). However, an interesting finding is that the LongFaith-PO synthesized with different base LLMs did not show significant performance differences in preference optimization. Even smaller model like Qwen2.5-7B-Instruct, are able to synthesize high-quality reasoning chains, with performance on some datasets matching or even surpassing GPT-4o mini. This highlights the robustness of the LongFaith pipeline, which is capable of modeling fine-grained preferences to synthesize high-quality instructions.
5 Conclusion
This paper addresses the challenge of questionable faithfulness in data synthesis approaches for long-context LLMs. We propose LongFaith, a novel pipeline synthesizing faithful long-context reasoning datasets through ground truth integration and citation-based reasoning prompts. Experiments demonstrate its effectiveness, with ablation studies confirming the adaptability of the LongFaith-SFT and LongFaith-PO datasets across diverse long-context scenarios.
Limitations
While LongFaith demonstrates significant improvements in long-context reasoning tasks, its scalability and generalization to other LLMs remain an open question. Our experiments focused on a single model, and thus, the performance of LongFaith on other general-purpose LLMs still needs further validation. Additionally, while the synthesized instruction sets with lengths of approximately 10,000 tokens successfully generalized to long-context reasoning tasks, future work will explore the extension of LongFaith to generate instructions with even longer contexts and evaluate the impact on model performance. Finally, LongFaith currently concentrates on reasoning tasks, and we plan to explore its generalization to other tasks such as summarization, dialogue generation, and others, to assess its broader applicability.
Appendix A An Example of Synthesized LongFaith-SFT and LongFaith-PO Datasets
Figure 6: An Example of synthesized LongFaith-SFT and LongFaith-PO datasets.
Appendix B Prompt Templates
We present the prompt templates that are used to synthesize the datasets. The core prompt template that generates long-context reasoning chains guided by ground truth using chain-of-citation is shown in Fig. 7. The samples are used in LongFaith-SFT dataset and are chosen as positive in LongFaith-PO dataset, since they are of the highest faithfulness. The other three prompt templates synthesize rejected samples for LongFaith-PO dataset, corresponding to (1) Misinformation due to lack of verification ( Fig. 8), (2) Reasoning without attribution (Fig. 9), and (3) Potential knowledge conflicts (Fig. 10).
Figure 7: An example of synthesized chosen reasoning chain. The current reasoning chain must arrive at a correct answer, and reasoning with proper citation proposes more faithfulness and interpretability. Therefore, LongFaith will choose it in supervised fine-tuning and preference optimization as positive sample.
Figure 8: An example of synthesized rejected reasoning chain. Misinformation due to lack of verification will cause more hallucination if we use current reasoning chain to fine-tune a LLM. Therefore, LongFaith will reject it in preference optimization.
Figure 9: An example of synthesized rejected reasoning chain. As it mentioned in previous work, lack of attribution will lead to much more interpretability and faithfulness, and response with citation is encouraged. Therefore, LongFaith will reject it in preference optimization.
Figure 10: An example of synthesized rejected reasoning chain. Information related to "Panama was colonized by Spain" is not mentioned in documents provided, which will cause knowledge conflicts if using current reasoning chain for fine-tuning. Therefore, LongFaith will reject it in preference optimization.
Appendix C Case Study of Metric "Hack" on SubEM comparing previous study with LongFaith
Figure 11: A case study that SubEM metric is "hacked" by previous study, which conduct more exploration with redundancy in response. LongFaith can arrive at the final correct answer with shorter response.
Appendix D Post-Training
In this section, we present two post-training algorithms—Supervised Fine-Tuning (SFT) and Preference Optimization (PO)—to better leverage synthetic data for efficiently enhancing model performance. Specifically, the model performs supervised fine-tuning on high-quality faithful outputs or is trained through reinforcement learning using synthetic preference pairs.
Supervised Fine-tuning on Faithful Outputs
We minimize the negative log-likelihood of the output as follows:
$\mathcal{L}_{\text{SFT}}$ $=-\frac{1}{|y|}\log\pi_{\theta}(y\;|\;x)$ (7)
$=-\frac{1}{|y|}\sum_{i=1}^{|y|}\log\pi_{\theta}(y_{i}\;|\;x,y_{<i})$
where $y$ denotes the high-quality faithful outputs, which are synthesized in Section 3.
Reinforcement Learning from Synthetic Preference
Additionally, we can leverage synthetic preference pairs for reinforcement learning (RL) to fine-tune the model toward generating faithful outputs while reducing the likelihood of low-scoring outputs. Standard RL algorithms for optimizing LLMs include Proximal Policy Optimization (PPO) Schulman et al. (2017), RLOO Ahmadian et al. (2024). However, these methods incur high computational costs. Recent approaches such as Direct Preference Optimization (DPO) Rafailov et al. (2024), Kahneman-Tversky Optimization (KTO) Ethayarajh et al. (2024), and Odds Ratio Preference Optimization (ORPO) Hong et al. (2024) have been proposed to mitigate both computational and data requirements. In this work, we adopt the ORPO algorithm, which achieves an optimal balance between computational efficiency and model performance.
ORPO introduces an odds ratio loss $\mathcal{L}_{OR}$ that minimizes the negative log odds ratio between preferred ("win" $y_{w}$) and dispreferred ("lose" $y_{l}$) outputs:
$\mathcal{L}_{\text{OR}}=-\log\sigma\left(\log\frac{\text{odds}_{\theta}(y_{w}|x)}{\text{odds}_{\theta}(y_{l}|x)}\right)$ (8)
where $\sigma$ denotes the sigmoid function and $\text{odds}_{\theta}(y|x)=\frac{\pi_{\theta}(y|x)}{1-\pi_{\theta}(y|x)}$ measures how much more likely $y$ is to be generated. The final objective of ORPO is to combine the SFT loss and the OR loss while controlling their relative importance through a hyperparameter $\beta$:
$\mathcal{L}_{\text{ORPO}}=\mathcal{L}_{\text{SFT}}+\beta\cdot\mathcal{L}_{\text{OR}}$ (9)
In this paper, the chosen output $y_{w}$ is synthesized by LongFaith through comprehensive consideration of Supporting Docs, Chain-of-Citation (CoC), and Ground Truth (GT), and is consequently assigned a high score. Conversely, the rejected output $y_{l}$ refers to synthesized outputs that lack at least one of these three critical elements (Supporting Docs, CoC, or GT), which are deemed low-scoring due to insufficient design considerations.
Appendix E Statistics of Main Experiments
MuSiQue #2-Hop #3-Hop #4-Hop
1K 0 512 512
2K 512 512 1024
4K 1024 2048 1024
8K 3072 4096 1024
Table 5: Statistics of train set for synthesis in different size sampled from MuSiQue Trivedi et al. (2022b).
Datasets #Count Avg. L. Max L.
Multi-Hop Reasoning
MuSiQue 500
2-Hop 254 10843.3 17560
3-Hop 154 11456.5 19225
4-Hop 92 11224.3 16756
2WikiMultiHopQA 500
2-Hop 395 4449.5 10631
4-Hop 105 4041.4 9365
HotpotQA 500
Bridge 412 6301.0 15702
Comparison 88 5777.6 11939
LongBench
Qasper (S) 200 24262.3 101636
MultiFieldQA-En (S) 150 29583.7 64751
MuSiQue (M) 200 69876.8 82338
2WikiMHQA (M) 200 30076.5 72971
HotpotQA (M) 200 57041.4 81815
Table 6: Statistics of test sets including three long-context multi-hop reasoning datasets sampled by Trivedi et al. (2022a) and five long-context QA datasets from LongBench Bai et al. (2023). Avg. L. and Max L. refer to the average length and max length of input prompts for test samples. (S) and (M) refer to Single-doc QA and Multi-doc QA in LongBench.
Datasets Instruction Output(Chosen) Rejected
LongAlpaca 52043.2 620.7 0
LongAlign 36307.2 1412.6 0
MuSiQue-Attribute 11395.0 343.7 0
LongMIT 280808.9 825.2 0
LongReward 72892.2 913.4 960.6
SeaLong 82248.6 1156.5 1139.1
LongFaith 11542.1 1029.6 896.7
Table 7: Average text length of baseline datasets and LongFaith in main experiments in Tab. 2. All of them has 2K examples.
Datasets MuSiQue 2WikiMultiHopQA HotpotQA Qasper MultiFieldQA Avg.L
LongAlpaca 365.62 372.97 319.60 657.34 511.25 445.36
LongAlign 493.56 349.65 371.77 651.76 623.15 497.98
MuSiQue-Attribute 99.61 168.74 164.24 317.75 252.95 200.66
LongMIT 138.03 159.16 116.40 194.41 196.43 160.89
LongReward-SFT 285.20 241.47 178.26 750.83 537.95 398.74
SeaLong-SFT 1091.54 776.01 926.29 1035.77 822.82 930.49
LongFaith-SFT 820.04 619.18 771.68 1056.55 941.13 841.72
LongReward-PO 219.90 253.41 179.44 616.11 460.40 345.85
SeaLong-PO 961.51 740.14 891.75 946.68 826.77 873.37
LongFaith-PO 831.17 669.76 786.77 1034.71 917.11 847.90
Table 8: Average length of model output in test sets trained on different synthesized instruction.
Appendix F Hyperparameters
Hyperparameters Value
# GPUs used 4
Learning rate 5e-5
Per-device batch size 1
Gradient accumulation steps 8
LoRA rank 32
LoRA alpha 64
LoRA dropout 0.1
ORPO beta 0.1
Warm-up ratio 0.1
Epochs 1
Precision bfloat16
Optimizer AdamW
Table 9: Hyperparameter settings of fine-tuning and preference optimization.